OpenStudy (anonymous):

Given that P(A) = 0.6 and P(B) = 0.3, and P(A and B) = 0.18, determine P(B|A).

OpenStudy (amistre64):

this the multilplication rule

OpenStudy (amistre64):

.18 = .6(.3) - b|a

OpenStudy (amistre64):

im going with 0

OpenStudy (anonymous):

=0.18/0.6

OpenStudy (amistre64):

p(a and b) = p(a).p(b) .18 = .6(.3) .... so p(b|a) =0 becasue the events are disjoint... if i recall correctly

OpenStudy (amistre64):

or..... maybe that p(b|a) = p(b); in that case.... .3 is right :)